# Enabling Ensemble Learning for Heterogeneous Large Language Models with Deep Parallel Collaboration

### v1

@@ 4.1 Experimental Setup
$\Delta$ +1.36 +1.35 +1.94 +2.14
Table 3: Results on knowledge-intensive QA Tasks. NQ refers to NaturalQuestion.
4.2 Results
Figure 3: Effect of the number of anchor words. The x-axis indicates the number of anchor words randomly sampled from the common words.
The main results are shown in Tab 4.1,4.1, and 4.1, from which we have drawn the following observations:
DeePEn achieves consistent improvements over the individual models across all benchmarks.
These results prove that our framework indeed enables collaboration between LLMs via aggregating the target semantics behind the output probability distributions. Specifically, the ensemble of top-2 models achieves the improvements of +1.5(NQ)$\sim$+3.22(PIQA) on test sets and +0.25(NQ)$\sim$+3.85(ARC-C) on validation sets in terms of accuracy. Through the ensemble of top-3 models, DeePEn gains a further improvement over ensembling top-2 models. As shown in Tab. 4.1, the improvement is increased from +1.79 to +2.99 after the joining of the third-ranked model on ARC-C.
Collaboration with models significantly worse could lead to degeneration.
As the results in Tab 4.1, the ensemble of top-3 models underperforms than the one of top-2 models on TrivialQA. This degeneration is reasonable due to the significant performance gap (-6.68) between the third-ranked model (InternLM-20B) and the second-ranked models (Mistral-7B) on this benchmark.
DeePEn shows excellent generalizability.
As we can see, Both PairRanker and GenFuser perform poorly on each benchmark, which is caused by the serious distribution shift from their training data to the test data. Instead, DeePEn works well under the training-free setting.
5 Analysis
Towards better understanding DeePEn, we conduct a series of analyses on the relative transformation and the algorithm of relative representation decoding.
@@ 5.1 Analysis on Relative Transformation
We demonstrate the impact of different numbers of anchor words through experiments with the top-2 ensemble models on the MMLU and ARC-C benchmarks. As shown in Fig. 3, we observe that an increased number of anchor words does not necessarily translate into improved performance for LLMs in downstream tasks. Fortunately, selecting the full set of common words as anchors could achieve stable promising performance.
@@
Methods MMLU-Dev TrivialQA-Dev
@@
Table 4: Ablation study of normalization on the relative representation matrix to the ensembling performance on the development sets. Baseline refers to as the best single model on each benchmark. DeePen refers the performance of ensembling top-2 models in the benchmark.
To demonstrate the importance of normalization on the relative representation matrix to the ensemble performance, we conduct an ablation analysis. The result is shown in Tab. 4, the ensemble struggles to achieve improvements due to the ineffective representation of outlier words, i.e., words distant to other words. The proportion of outlier words can be derived from the distribution of distance to nearest neighbor words, which is illustrated in Fig. 4. As illustrated, a remarkable proportion ($>30\%$) of words are distant from other words, i.e., cosine similarity to its nearest neighbor word is less than 0.3. Through the normalization operation, the target semantics that intend to emit outlier words could be prevented from becoming zero vectors by relative transformation.
@@ 5.2 Analysis of Reverse Transformation
As described in §3.4, the relative representation of the aggregated target semantic is transformed back to the absolute space of the main model via a search-based algorithm. There are two important factors (relative ensemble learning rate and number of search steps) in this reverse transformation, which are comprehensively analyzed, respectively.
@@
Baseline TrivQA NQ ARC-C MMLU
TrivialQA 72.74 72.90 72.13 69.35 70.70
NQ 21.56 21.56 21.88 21.81 21.62
ARC-C 59.32 69.32 71.97 73.76 73.26
MMLU 55.10 59.98 61.02 61.86 61.42
Table 6: Cross-distribution validation of relative ensemble learning rate ($\eta$). We report the performance of ensembling LLaMA2-13B and Mistral-7B. Each row indicates the test set used to evaluate performance. Each column indicates the development set used to search the optimal value of RELR.
Analysis of relative ensemble learning rates.
As shown in Tab. 5, the performance of DeePEn is sensitive to the value of relative ensemble learning rate ($\eta$), which is abbreviated by RELR. This observation motivates us to measure the generality of this hyperparameter. Specifically, we illustrate the cross-distribution performance of the searched optimal value of $\eta$ in Tab. 6. As observed, the optimal value of RELR varies across different datasets, which suggests that the inverse transformation from relative space to absolute space requires adaptive mapping modes.
@@
To give a deep view of the dynamics of the inverse transformation in DeePEn, we report the performance change along with different numbers of relative ensemble learning steps ($T$). Besides, the dynamics of loss of relative ensemble learning ($\ell$ in Eq. 8)is also reported. As shown in Fig. 5, on the one hand, more steps of relative ensemble learning significantly lead to lower losses. However, the loss is hard to reach zero, i.e., under-fitting. On the other hand, increasing the number of steps of relative ensemble learning will cause the performance to increase first and then decrease. The reason behind the performance drop could be that in the early stage of optimization, the focus of optimization is on updating the high-probability tokens. In the later stage of optimization, since the probabilities of all words will be adjusted equally, the low-probability tokens will be interfered with, thus affecting the performance. Therefore, it is recommended to set a modest value of step number (e.g., $T=5$).
Choice of main model.
@@
In the process of inverse transformation, DeePEn maps the relative aggregated representation to the absolute space of the main model. Ideally, we expected the results of inverse transformation to keep invariant with the choice of main model. However, this objective is hard to achieve due to the underfitting observed in the search process. Therefore, we illustrate the performance gap of choosing different main models in Tab. 7. As the results shown on ARC-C, changing the main model from the first-ranked Mistral-7B to the second-rank Yi-6B, the ensemble performance is decreased slightly from $77.73$ to $77.55$. Interestingly, changing the main model from the rank-1 Yi-6B to the rank-2 Mistral-7B on MMLU, the performance is actually improved from $63.63$ to $64.46$, which indicates that Mistral-7B benefits more than Yi-6B from collaboration. Even so, choosing different main models does not significantly affects the ensemble performance.
5.3 Comparison to Vanilla Prediction Average
@@
To compare our DeePEn with vanilla prediction average, we conduct an experiment for ensembling two LLMs with the same vocabulary and comparable performance on MMLU, i.e., LLaMA2-7B and LLaMA1-13B. As shown in Tab. 8, the performance of DeePEn is comparable, even better than, that of the vanilla prediction average. Theoretically, the performance of the vanilla prediction average is the performance upper-bound of DeePEn. The reason that DeePEn could excel over the vanilla one on MMLU is the under-fitting in the inverse transformation process, which leads to the weights to aggregate the target semantics of different models not being a uniform distribution (i.e., $(0.5,0.5)$). For example, in Tab. 8, the weights for LLaMA1 and LLaMA2 could be $(0.6,0.4)$, where the weight of the main model is larger than the other model.
6 Related Work
Existing work of LLM ensemble could be divided into selection-based and fusion-based ensemble.
Selection-based ensemble.
Rerank is an intuitive solution to utilize multi-model strengths. Specifically, Jiang et al. (2023b) takes the first step towards ensembling LLMs, proposing PairRanker for pairwise comparison on candidate outputs and achieving improvements on the self-constructed instruction dataset. To overcome the huge computation costs of multi-LLM inference, several works have explored to train a router to predict the best-performing model out of a fixed set of LLMs for the given input (Wang et al., 2024; Shnitzer et al., 2024; Lu et al., 2023).
Fusion-based ensemble.
Towards a synergy between LLMs, Jiang et al. (2023b) propose GenFuser, trained to generate an improved response to capitalize on the strengths of multiple candidates. Furthermore, Du et al. (2023) design a debate-like prompt strategy to ask the LLM to refine its answer after reading peer models’ answers. However, this prompt-based method incurs substantial computation costs and inference latency for LLM communication. Specifically, considering the ensemble of $N$ LLMs, the communication leads to the encoding of $N\times L$ tokens and decoding of $L$ tokens, where $L$ denotes the average length of answers.
Different from the training-dependent ensemble methods which pose a great challenge to the generalizability of the trained reward model or fusion model, our DeePEn is training-free, making it more general. Compared to debate-like prompt strategy, DeePEn has higher computation efficiency and achieves more stable improvements.We place the comparison experiments with Debate in Appendix A.1.
A contemporaneous work to us is FuseLLM, which propose a method of vocabulary mapping to enable knowledge distillation between heterogeneous LLMs. Even though FuseLLM works for knowledge distillation, it hardly achieve improvement under the training-free setting of ensemble learning, as the experimental results shown in A.2.
7 Conclusion
In this work, we propose a training-free LLMs collaboration framework DeePEn, which averages the probability distributions of models with heterogeneous vocabularies. To the best of our knowledge, we are the first to effectively enable the aggregation of heterogeneous probability distributions. We believe our research can motivate further research on the LLMs collaboration, model reuse, and knowledge distillation. In the future, we are planning to experiment with DeePEn on the ensemble between LLMs and generative expert language models (e.g., machine translation models) due to their complementary knowledge.
8 Limitations
Different from previous LLM ensemble methods, our DeePEn is uniquely designed for the ensemble of white-box LLMs due to necessitating authorization of the output probability distribution. Besides, DeePEn assigns equal weights to each individual model, which fails to deal with the interference caused by the poor-performing model. This problem motivates us to introduce ensemble strategies like stacking technology (Wolpert, 1992).
Appendix A Comparison with More Baselines
Models ARC-C-Test
Acc (w/o. CoT) Acc (w. CoT) Communication Cost Inference Latency
Mistral-7B 74.10 67.86 0 1x
Yi-6B 73.33 70.60 0 1x
Debate 73.25 (-0.85) 72.20 (+1.4) 14B×$\times L$ 2x
FuseLLM 74.10 (+0.00) 0.20B×$\times L$ 1.0x
DeePEn 75.89 (+1.79) 0.85B×$\times L$ 1.25x
Table 9: Comparison with previous debate-like prompt strategy and FuseLLM in terms of performance (Accuracy), computation cost for communication, and inference latency. $L$ represents the average length of generated response.
A.1 DeePEn vs. Debate
We compare our DeePEn with the debate-like prompt strategy (Du et al., 2023) on the ARC-C benchmark in Tab. 9. As we can see, Debate works only on the CoTWei et al. (2022) setting while underperforming on the setting without CoT. Note that DeePEn achieve a larger improvement (+1.79) over a stronger baseline (74.10), which indicate the superiority. Besides, DeePEn incurs less computation cost than Debate. Specifically, the cost of DeePEn is estimated as $(N+2T-1)\times|V|\times|A|$, and the cost of Debate is estimated as $N\times L\times|\theta|$. Last but not least, DeePEn leads to less inference latency than Debate.
A.2 DeePEn vs. FuseLLM
To achieve a fair comparison to FuseLLM, we re-implement their vocabulary mapping method with the released code under the ensemeble learning setting. Specifically, on ARC-C, the probability distribution outputted by Yi-6B is mapped in the vocabulary of Mistral-7B. Then, we averaged the mapped probability distribution and the probability distribution outputted by Mistral-7B. Finally, the averaged distribution is used to determine the emitted token. As the results in Tab. 9, FuseLLM hardly achieve improvements under the setting of ensemble learning. The most essential difference is that FuseLLM uses the literal similarity to enable the token alignment while DeePEn leverages the relative representation to enable the probability distribution aggregation.

### v2

@@ 5.1 Results of Ensembling Diverse Models
$\Delta$ +1.60 +3.22
Figure 4: Ensemble learning of the dense large language model LLaMA2-70B and the sparse MoE model Mixtral-8$\times$7B.
[…]
LLaMA2-13B 30.60 42.27 30.83 39.99
NLLB-600M 32.30 41.49 31.91 42.39
[…]
$\Delta$ +1.04 +1.43 +1.04 + 0.45
Figure 5: Ensemble learning of the generalist model LLaMA2 and the specialist translator model NLLB on the translation benchmark Flores-200.
Ensemble of the dense model and the sparse model.
We first evaluate our method on the ensemble learning of the dense model and the sparse MoE model on the challenge reasoning tasks. Specifically, we use the widely-used large-scale dense model LLaMA2-70B [29] and the popular sparse MoE model Mixtral-8$\times$7B [14] as the base models. As the results shown in Tab. 5.1, our DeePEn achieves improvements of +1.60 and +3.22 on the GSM8K and PIQA datasets, even though the base models have achieved a high level of performance.
Ensemble of the generalist model and the specialist model.
To investigate the effectiveness of DeePEn on the ensemble of the generalist model and the specialist model for the specific task, we conduct experiments on the machine translation task using the ensemble of the large language model LLaMA2 and the machine translation model NLLB [27], which is a well-known open-source multilingual translator. We adopt the widely-used machine translation benchmark Flores-200222https://github.com/facebookresearch/flores. As the results in Tab. 5.1 illustrated, DeePEn achieves better translation performance leveraging the diverse translation knowledge in the generalist LLM and the specialist translator.
Ensemble of models with different performance gaps.
To assess the stability of DeePEn regarding to the performance gap of base models, we conduct an experiment on the ensemble of model pairs with increasing performance gaps. As the result demonstrated in Tab. 8, the performance of ensemble learning between a well-performing model (the rank-first model)with a worse-performing model could achieve improvements or slightly lag behind the well-performing model.
@@ 5.2 Analysis on Relative Transformation
We demonstrate the impact of different numbers of anchor words through experiments with the top-2 ensemble models on the MMLU and ARC-C datasets. As shown in Fig. 6, an increased number of anchor words can improve performance for LLMs in downstream tasks, and selecting the full set of common words as anchors provides better performance.
@@
Figure 6: Effect of the number of anchor words. The x-axis indicates the number of anchor words randomly sampled from the common words for 4 times.
Methods MMLU-Dev TriviaQA-Dev
@@
Figure 7: Ablation study of normalization on the relative representation matrix to the ensembling performance on the development sets. Baseline refers to as the best single model on each benchmark. DeePen refers the performance of ensembling top-2 models in the benchmark.
Figure 8: 2-model ensemble of the top-1 model (LLaMA2-13B) with different models on the NQ benchmark, respectively.
To demonstrate the importance of normalization on the relative representation matrix to the ensemble performance (§3.2), we conduct an ablation analysis. The result is shown in Tab. 8, the ensemble struggles to achieve improvements due to the ineffective representation of outlier words, i.e., words distant to other words. The proportion of outlier words can be derived from the distribution of distance to nearest neighbor words, which is illustrated in Fig. 11. As illustrated, a remarkable proportion ($>30\%$) of words are distant from other words, i.e., cosine similarity to its nearest neighbor word is less than 0.3. Through the normalization operation, the output semantics that intend to emit outlier words could be prevented from becoming zero vectors by relative transformation.
@@ 5.3 Analysis of Reverse Transformation
To better understand the reverse transformation process (§3.4) transforming the relative representation back to the absolute space of the main model, we further analyze each component of this process.
Analysis of relative ensemble learning rates.
As shown in Tab. 2, the performance of DeePEn is sensitive to the value of relative ensemble learning rate ($\eta$), which is abbreviated by RELR. This observation motivates us to measure the generality of this hyperparameter. Specifically, we illustrate the cross-distribution performance of the searched optimal value of $\eta$ in Tab. 6. As observed, the optimal value of RELR varies across different datasets, which suggests that the inverse transformation from relative space to absolute space requires adaptive mapping schemes.
@@
To give a deep view of the dynamics of the inverse transformation in DeePEn, we report the performance change along with different numbers of relative ensemble learning steps ($T$). Besides, the dynamics of loss of relative ensemble learning ($\eta$ in Eq. 6)is also reported. As shown in Fig. 12, on the one hand, more steps of relative ensemble learning significantly lead to lower losses. However, the loss is hard to reach zero, i.e., under-fitting. On the other hand, increasing the number of steps of relative ensemble learning will cause the performance to increase first and then decrease. The reason behind the performance drop could be that in the early stage of optimization, the focus of optimization is on updating the high-probability tokens. In the later stage of optimization, since the probabilities of all words will be adjusted equally, the high-probability tokens will be interfered with the high-probability ones, thus affecting the performance. Therefore, it is recommended to set a modest value of step number (e.g., $T=5$).
6 Related Work
Selection-based ensemble.
Rerank is an intuitive solution to utilize multi-model strengths. Jiang et al. [15] take the first step towards LLM ensemble, training a reward model PairRanker for pairwise comparison on candidate outputs. To overcome the huge computation costs of multi-LLM inference, several works have explored to train a router to predict the best-performing model out of a fixed set of LLMs for the given input [31, 25, 19].
Fusion-based ensemble.
Towards a synergy between LLMs, Jiang et al. [15] propose GenFuser, trained to combine multiple candidate answers. Different from these training-dependent ensemble methods which pose a great challenge to the generalizability of the reward model or fusion model, our DeePEn is completely training-free, making it more general. Similar to our method, MinED also aims to tackle the vocabulary discrepancy via aligning the tokens in different vocabularies based on edit distance [30, 9]. Unfortunately, this textual similarity-based method exhibits unstable performance and produces abnormal text for LLM ensemble (Tab. 10).
There are several contemporaneous works related to our work. Xu et al. [33] propose EVA to tackle vocabulary discrepancy by learning token alignment between different vocabularies with the assistance of overlapping tokens. Our DeePEn eliminates this training process via directly aligning tokens with the relative representation (more discussion is illustrated in §B). Mavromatis et al. [20] explore adaptive collaboration weights at test time by harnessing the perplexity on the input prompt. We emphasize that this work is complementary to our work.
7 Conclusion
In this work, we propose a training-free LLM ensembling framework DeePEn, which addresses the vocabulary discrepancy when fusing the probability distributions of heterogeneous LLMs. Experimental results on six widely-used benchmarks demonstrate that DeePEn exhibits more stable performance than baseline methods and has complementary strengths with other ensemble methods such as Voting. We believe our work can inspire further research on the LLMs collaboration, model reuse, and knowledge distillation. In the future, we aim to explore more effective adaptive collaboration schemes to leverage the complementary strengths between different LLMs.
Appendix A Statistics of Common Tokens across different LLMs
We count the number of common tokens shared among different LLM vocabularies and present the results in Fig. 9. It is observed that a large number of common words (>20k) exist across the different vocabularies. We also count the number of common tokens in all six LLMs and find that there are a total of 18k common tokens, enabling DeePEn to be applied to the ensemble learning of a large number of models.
Figure 9: Statistics of common words across different vocabularies.
Appendix B Details of Baselines
LLM-Blender.
(1) the selection-based ensemble method PairRanker Jiang et al. [15], which is a reward model to score each response of LLMs and (2) the fusion-based ensemble method GenFuser Jiang et al. [15], which is a generative model to fuse multiple candidate responses. Both models are trained on the constructed instruction tuning dataset MixInstruct. In our experiments, as GenFuser struggles to generate responses following the expected format, we only adopt PairRanker.
Voting.
For tasks with outputs limited to a fixed set (i.e., MMLU, ARC-C, PIQA, GSM8K benchmarks), we adopt the Voting method on the ensemble learning of more than 2 models. Concretely, we count each candidate answer’s occurrences and select the most frequent as the final output. In the event of a tie, the main model’s answer is used as the final output.
MBR.
For generation tasks, we implement the MBR [8, 17] method, which selects the answer with the highest lexical similarity to other candidate answers. To measure this similarity, we experimented with the edit distance and chrF [^3] https://github.com/mjpost/sacrebleu metrics, ultimately choosing chrF due to its superior performance.
MinED.
To bridge the gap between different vocabularies in LLM ensemble, MinED apply the Minimum Edit Distance (MinED) approach to align tokens across different vocabularies, e.g., "get" to "gets". However, this textual similarity-based mapping method could disturb the text generation process and produce incomplete words.
EVA.
Recently, Xu et al. [33] propose EVA to tackle the vocabulary discrepancy by learning mappings between the vocabularies of different LLMs with the assistance of overlapping tokens. We have tried to re-implement their method with the released code. However, we encounter a technical problem in that EVA only the supports the ensemble learning between LLMs with the same embedding dimension. This is caused by the limitation of tool of vecmap [^4] https://github.com/artetxem/vecmap, which is used to learn the token alignment.
Figure 10: Analysis of the generation process of MinED. To illustrate the problematic generation process of MinED, we list the top-10 high-probability tokens in the probability distribution of the assistant model and their aligned token.
Appendix C Additional Experiments
C.1 Choice of main model.
@@ Appendix C Additional Experiments
In the process of inverse transformation, DeePEn maps the relative aggregated representation to the absolute space of the main model. Ideally, we expected the results of inverse transformation to keep invariant with the choice of the main model. However, this objective is hard to achieve due to the underfitting observed in the search process. Therefore, we illustrate the performance variance of choosing different main models in Tab. 3. As the results shown on ARC-C, changing the main model from the first-ranked Mistral-7B to the second-rank Yi-6B, the ensemble performance is decreased slightly from $77.73$ to $77.55$. Interestingly, changing the main model from the rank-1 Yi-6B to the rank-2 Mistral-7B on MMLU, the performance is actually improved from $63.63$ to $64.46$, which indicates that Mistral-7B benefits more than Yi-6B from collaboration. Even so, choosing different main models does not significantly affects the ensemble performance.
C.2 Comparison to Vanilla Prediction Average
@@
To compare our DeePEn with vanilla prediction average, we conduct an experiment for ensembling two LLMs with the same vocabulary and comparable performance on MMLU, i.e., LLaMA2-7B and LLaMA1-13B. As shown in Tab. 4, the performance of DeePEn is comparable, even better than, that of the vanilla prediction average. Theoretically, the performance of the vanilla prediction average is the performance upper-bound of DeePEn. The reason that DeePEn could excel over the vanilla one on MMLU is the under-fitting in the inverse transformation process, which leads to the weights to aggregate the output semantics of different models not being a uniform distribution (i.e., $(0.5,0.5)$). For example, in Tab. 4, the weights for LLaMA1 and LLaMA2 could be $(0.6,0.4)$, where the weight of the main model is larger than the other model.
C.3 Latency Analysis
To accomplish the fusion of heterogeneous distributions, DeePEn first maps the distributions into the relative space and adopts the search-based inverse transformation to map the aggregated relative representation back to the main model’s probability distribution, which incurs an extra latency. This latency is mainly caused by the inverse transformation process, which requires $T$-round search. To demonstrate this latency, we report the token-level inference latency of ensembling two LLMs (Mixtral-8$\times$7b and LLaMA2-70B). This experiment is conducted on 8 A100 GPUs. All of our experiments can be re-implemented on 8 A100 GPUs. As shown in Tab. 5, DeePEn causes +17% token-level inference latency. However, in practice, this latency could be greatly decreased since all individual models intend to emit the same token in 90% decoding steps. In these steps, we could skip the fusion process and use the consistently agreed token as the next token. In total, DeePEn actually incurs less than 2% sentence-level inference latency.
Baseline $T=1$ $T=3$ $T=5$ $T=10$
Inference Latency 0.19s 0.20s 0.21s 0.22s 0.24s
Relative Change 0% +7% +11% +17% +29%
Table 5: Inference Latency of DeePEn with different search steps $T$.
Baseline TrivaQA NQ ARC-C MMLU
TriviaQA 73.42 75.9 75.41 75.56 75.44
NQ 29.11 30.55 30.65 30.42 30.69
ARC-C 60.29 69.32 72.31 74.19 73.76
MMLU 54.06 59.97 61.04 61.94 61.42
Table 6: Cross-distribution validation of relative ensemble learning rate ($\eta$). We report the performance of ensembling LLaMA2-13B and Mistral-7B. Each row indicates the test set used to evaluate performance. Each column indicates the development set used to search the optimal value of $\eta$.
Appendix D Limitations
As illustrated in Tab. 4.1, collaboration with more LLMs can sometimes lead to a performance drop caused by interference from lower-performing models. This issue limits the ensemble performance of our current method, even though we have explored setting different collaboration weights for each model on each benchmark (DeePEn-Adapt). An ideal solution would be to set adaptive collaboration weights at the sample level, or even the token level, for each LLM, which remains a significant challenge. Despite this, our work represents an important step towards the distribution fusion of LLMs.